### v1

# Shuffle Mamba: State Space Models with Random Shuffle for Multi-Modal Image Fusion

Abstract
Multi-modal image fusion integrates complementary information from different modalities to produce enhanced and informative images. Although State-Space Models, such as Mamba, are proficient in long-range modeling with linear complexity, most Mamba-based approaches use fixed scanning strategies, which can introduce biased prior information. To mitigate this issue, we propose a novel Bayesian-inspired scanning strategy called Random Shuffle, supplemented by an theoretically-feasible inverse shuffle to maintain information coordination invariance, aiming to eliminate biases associated with fixed sequence scanning. Based on this transformation pair, we customized the Shuffle Mamba Framework, penetrating modality-aware information representation and cross-modality information interaction across spatial and channel axes to ensure robust interaction and an unbiased global receptive field for multi-modal image fusion. Furthermore, we develop a testing methodology based on Monte-Carlo averaging to ensure the model’s output aligns more closely with expected results. Extensive experiments across multiple multi-modal image fusion tasks demonstrate the effectiveness of our proposed method, yielding excellent fusion quality over state-of-the-art alternatives. Code will be available upon acceptance.
Introduction
Multi-modal image fusion, a fundamental task in computer vision, involves extracting and integrating valuable information from images of the same scene captured by different imaging modalities. This task aims to create a single composite image with a more comprehensive and informative representation, with typical applications including pan-sharpening and medical image fusion (MIF). In the context of pan-sharpening, satellites are limited by sensors, which can only capture low-resolution multi-spectral (LRMS) and panchromatic (PAN) images. Specifically, PAN images offer superior spatial details but limited spectral resolution, while MS images provide abundant spectral resolution but lack spatial clarity. Through integrating the complementary information from both MS and PAN images into a composite representation, we can achieve an effective balance between spatial and spectral resolution. Analogously, in the realm of MIF, various imaging technologies capture distinct types of information. For instance, Computed Tomography (CT) images deliver detailed insights into bones and high-density tissues, while Magnetic Resonance Imaging (MRI) offers higher-resolution images with rich soft tissue details. In virtue of this complementary information from various modalities, MIF can overcome the limitations of single-modality images, resulting in a more comprehensive and detailed representation for modern medical diagnosis.
Figure 1: Visualization of Effective Receptive Fields (ERFs) for Conv, Self-Attention and our Method. A larger ERF is indicated by a more extensively distributed dark area.
In recent years, the prosperous advancement of deep neural networks (DNNs) has led to the development of numerous DNN-based multi-modal image fusion methods. In pan-sharpening, the pioneering work PNN (Masi et al. 2016) employed a simple three-layer neural network to achieve remarkable results which was previously deemed unattainable, highlighting the superior learning capabilities of deep learning. After that,increasingly complex and deeper architectures have emerged (Zhou et al. 2022b; Yang et al. 2023), delivering excellent visual performance. Despite these advancements, existing multi-modal fusion methods face common limitations. Convolutional neural network (CNN)-based approaches often struggle to establish global receptive fields. While transformers address this issue through self-attention mechanisms, they introduce significant challenges related to quadratic computational complexity. Nowadays, structured state-space models have gained considerable attention for their computational efficiency and principled ability to model long-range dependencies. However, their selective scanning mechanism can introduce biased priors when processing 2D images. To overcome these challenges, it is reasonable for us to design a novel sequential scanning method with its application framework.
Our Motivation. Global modeling capabilities are essential in image restoration tasks because one of the core aspects of image restoration is to find useful information within the image to compensate for the missing data in the current patch. For a long time, CNN and ViT have been dominant architectures in computer vision, each with distinct advantages and limitations. CNNs are constrained by local receptive fields, which hinder their ability to model long-range dependencies. In contrast, ViT uses the self-attention mechanism to access a global receptive field but are burdened by quadratic computational complexity. Recently, structured state-space models have demonstrated enhanced capabilities in capturing long-term dependencies in sequence data while maintaining linear time complexity. Notably, Mamba has achieved significant strides in reducing inference latency and improving overall performance through selective state spaces and hardware-aware algorithms. With the introduction of vmamba (Liu et al. 2024) and Vision Mamba (Zhu et al. 2024), there has been a growing interest in applying state-space models to visual tasks.
However, most current Mamba-based methods employ unidirectional SSMs, which endure certain limitations in their receptive field range due to their scanning approaches. Specifically, the receptive field is large for the initial portion of input patches but becomes significantly smaller for the latter portion, potentially compromising the model’s global modeling capabilities. Moreover, dissimilar to the dependencies between token orders in language modeling, the non-causal nature of 2-D spatial information in images presents a considerable challenge for simple sequential scanning methods. Traditional strategies, such as flattening image patches and scanning them sequentially, may introduce biased local 2D dependencies, thereby undermining the model’s ability to effectively identify spatial relationships.
To address these challenges, we propose a new sequence scanning method called Random Shuffle Scanning. Figure 1 shows the visualization of the effective receptive field (ERF) of three methods. Compared to the Conv method, the Self-Attention method’s advantage lies in achieving a global receptive field. However, our method provides an overall unbiased global receptive field, enabling the network to discard fixed local priors and focus more effectively on what needs to be learned. For the sequential input image patches, we first apply position encoding, followed by a random shuffle of the patches before they are processed by the Mamba block for long-range dependency modeling. The random shuffle approach eliminates the preconception between local and global 2D dependencies in mathematical expectation, enabling the model to access an unbiased prior and ultimately establish a more consistent and global receptive field. Recognizing that shuffling the spatial position of the image patches may disrupt semantic information, we implement a corresponding inverse transformation to accurately restore the sequence order of the input patches after passing through the Mamba block. This information-preserving transformation pair underpins the key components of our model: Random Mamba Block, Random Channel Interactive Mamba Block, and Random Modal Interactive Mamba Block. Inspired by Dropout, we use Monte Carlo averaging to approximate the expected output, ensuring that the actual output at test time closely matches the anticipated results.
Our contributions can be summarized as follows: 1) We design the Shuffle Mamba framework, where the random shuffle operation in the key component provides an expected unbiased global receptive field without increasing any parameters. 2) We develop a specific strategy for training and testing this framework. During training, each input is independently scanned using a random shuffle operation. During testing, we use Monte Carlo averaging to estimate the output of each Mamba block. 3) Extensive experiments on two prominent multi-modal image fusion tasks demonstrate that our method accomplishes excellent performance in both quantitative assessments and visual quality.
Related Works
State Space Model
The State Space Model (SSM), originating from control theory, has found extensive applications in deep learning due to its remarkable ability to model long-range dependencies. Initially, the S4 (Gu, Goel, and Ré 2021) model introduced the concept of the SSM, effectively reducing the computational and memory requirements associated with state representation while enabling global information modeling. Building upon S4, the S5 (Smith, Warrington, and Linderman 2022) model features MIMO structure and efficient parallel scanning strategy to enhance performance without significantly increasing computational demands. Furthermore, the H3 (Mehta et al. 2022) model further refines these methods, achieving competitive performance and efficiency comparable to the transformer in language modeling tasks.
Figure 2: The architecture of the proposed Shuffle Mamba framework.
Recently, Mamba (Gu and Dao 2023) has significantly improved inference speed and performance metrics through selective state spaces and hardware-aware algorithms. The introduction of Vmamba (Liu et al. 2024) and Vision Mamba (Zhu et al. 2024) has brought attention to the application of SSM in vision tasks. However, most existing SSM-based vision models (Zhu et al. 2024; Ma, Li, and Wang 2024; Zheng and Wu 2024) employ a fixed scanning strategy, which may introduce preconceptions, particularly in low-level vision tasks (Xiao et al. 2023). Specifically, this fixed order of choosing image patches can cause the model to gradually dismiss previous input sequences while processing the current patch, thereby compromising its capability to model global information. To alleviate this challenge, Vmamba (Liu et al. 2024) introduced the CSM, which scans image pixels from various directions such as top-left, bottom-right, top-right, and bottom-left. Building on this idea, RSM (Zhao et al. 2024) devised the OSSM to flatten image patches into sequences in eight directions, enhancing the network’s capability to capture and model large-scale spatial features. Additionally, LocalMamba (Huang et al. 2024) applies the windowed selective scan to ensure a harmonious integration of global and local visual cues. RSMamba (Chen et al. 2024) incorporates dynamic multi-path activation mechanisms to model non-causal data. However, these methods do not simultaneously consider the integrity of the image’s structure and the randomness of the pixels during scanning processing, and are essentially still special fixed strategies. To address this limitation, we introduced the Shuffle Mamba framework, which includes random shuffle and inverse operations to obtain a global receptive field without bias in expectations.
Method
Preliminaries
Inspired by continuous linear time-invariant (LTI) systems, SSMs exploit an implicit latent state $h(t)\in{\mathbf{R}^{N}}$to map a 1-D sequence $x(t)\in\mathbf{R}$ to $y(t)\in\mathbf{R}$. Specifically, SSMs can be mathematically expressed as an ordinary differential equation (ODE):
${h}^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)=\mathbf{C}h(t).$ (2)
Where $\mathbf{A}\in{\mathbf{R}^{N\times N}}$ is the evolution matrix, while $\mathbf{B}\in{\mathbf{R}^{N\times 1}}$ and $\mathbf{C}\in{\mathbf{R}^{1\times N}}$ serve as the projection parameters. However, solving these differential equations in a deep learning context can be challenging. The S4 and Mamba models propose introducing a timescale parameter $\mathbf{\Delta}$ to convert continuous parameters $\mathbf{A}$, $\mathbf{B}$ into their discrete counterparts $\bar{\mathbf{A}}$, $\bar{\mathbf{B}}$:
${{h}_{\text{t}}}=\bar{\mathbf{A}}{{h}_{t-1}}+\bar{\mathbf{B}}{{x}_{t}},$ (3)
${{y}_{t}}=\mathbf{C}{{h}_{t}},$ (4)
$\bar{\mathbf{A}}=\mathbf{\exp}\mathbf{(\Delta A)},$ (5)
$\bar{\mathbf{B}}={{\mathbf{(\Delta A)}}^{-1}}(\exp\mathbf{(\Delta A)-I)}\cdot\mathbf{\Delta B}.$ (6)
Finally, the output of the system can be attained through global convolution:
$\bar{\mathbf{K}}=(\mathbf{C\bar{B}},\mathbf{C\bar{A}\bar{B}},...,\mathbf{C{{\bar{A}}^{L-1}}\bar{B}})$ (7)
where $\mathbf{L}$ represents the length of the sequence $x$, $\bar{\mathbf{K}}\in\mathbf{{R}^{L}}$ is a structured convolution kernel.
Network Architecture
The proposed Shuffle Mamba framework consists of three functional components: the Random Mamba block (RM block), the Random Channel Interactive Mamba block (RCIM block), and the Random Modal Interactive Mamba block (RMIM block). The overall workflow is illustrated in Figure 2. Assuming that the input images with different modalities are ${\mathbf{M}_{up}}$ and ${\mathbf{P}_{in}}$, where ${\mathbf{M}_{up}}$ serve as the upsampled ${\mathbf{M}_{in}}$, we first use convolutional layers to project the images into the feature space. Given the limited receptive field of the convolutional layers, which makes capturing global features challenging, we perform patch embedding and send the resulting patches to the RM block for global feature extraction. This process yields the global modality-specific features $\mathbf{F}_{n}^{m}$ and $\mathbf{F}_{n}^{p}$, where $PE=PatchEmbed$:
$\mathbf{F}_{0}^{m},\mathbf{F}_{0}^{p}=PE(\phi({\mathbf{M}_{up}})),PE(\phi({\mathbf{P}_{in}}))$ (8)
$\mathbf{F}_{n}^{m}={{\psi}_{1n}}\cdot\cdot\cdot({{\psi}_{11}}({{\psi}_{10}}(\mathbf{F}_{0}^{m}))$ (9)
$\mathbf{F}_{n}^{p}={{\psi}_{2n}}\cdot\cdot\cdot({{\psi}_{21}}({{\psi}_{20}}(\mathbf{F}_{0}^{p}))$ (10)
The global modality-specific features are then sent to the RCIM block for simple channel information exchange, without introducing additional parameters. The exchanged features continue to be processed by their respective RM block to obtain $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$. Next, we use the RMIM block to attain the $\mathbf{F}_{k+j}^{m}$ through the deep fusion of modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$. Thus, the final fused image ${\mathbf{H}_{out}}$ can be accessed by reshaping and residual connection:
$\mathbf{F}_{k+j}^{m}={{\theta}_{j}}\cdot\cdot\cdot({{\theta}_{1}}({{\theta}_{0}}(\mathbf{F}_{k}^{m},\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p})$ (11)
${\mathbf{H}_{out}}=\phi(reshape(\mathbf{F}_{k+j}^{m}))+{\mathbf{M}_{up}}$ (12)
Figure 3: The Random Shuffle Scanning for training.
Figure 4: The Monte-Carlo averaging for testing.
Key Components
Random Shuffle Scanning.
Mamba was originally designed to adapt to the modeling of language sequences. We devised a Random Shuffle Scanning approach to access unbiased local and global dependencies in 2D image while ensuring a global receptive field similar to self-attention. As shown in Figure 3, for sequentially input 2D image patches, we first apply depth-wise convolution for position modeling. The image patches are then randomly shuffled and sent to the Mamba block for long-range dependency modeling. This strategy allows the Mamba block to simulate interactions between adjacent patches with equal probability, enabling the network to learn and model from an unbiased prior effectively. Additionally, since the relative positions of the patches are crucial for reconstructing semantic information, the output image must be accurately aligned with the input based on the inverse shuffle. For this reason, the random shuffle and its inverse operation constitute an information-lossless transformation pair.
Random Mamba Block.
Based on this shuffle-inverse pair, we designed the Random Mamba Block. First, layer normalization is performed on the input feature ${\mathbf{F}_{l-1}}$ to obtain ${{\mathbf{F}^{\prime}}_{l-1}}$ , which is then projected into $\mathbf{x}$ and $\mathbf{z}$ using random shuffle and multi-layer perceptrons (MLPs). In the first branch, $\mathbf{x}$ passes through 1-D convolution layers with SiLU activation to produce $\mathbf{x}^{\prime}$. The SSM is then used to calculate the output $\mathbf{y}$. In the other branch, $\mathbf{z}$ is sent to the activation function to generate the gating for $\mathbf{y}$. Finally, we apply the inverse shuffle and residual connection to get the final output sequence ${\mathbf{F}_{l}}$.
Random Channel Interactive Mamba Block.
In the RCIM block, we refer to the approach from (He et al. 2024a) to achieve lightweight feature interaction between different modalities. We use the split operation to divide modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$ into two halves based on the channel dimension, followed by complementary splicing. The exchanged features are then sent to their respective RM Blocks for processing. By repeating these steps, the global modality-specific features are initially fused.
Random Modal Interactive Mamba block.
Motivated by cross-attention, we designed the RMIM block for processing multi-modal image information. In this approach, we project the shuffled sequence features into a shared space and use a gating mechanism to learn complementary information under an unbiased prior, thereby reducing the interference of redundant features on the fusion results. We employ a process similar to RM Block to generate ${\mathbf{y}_{m}}$ and ${\mathbf{y}_{p}}$, and use the input $\mathbf{F}_{l-1}^{m}$ to generate the gating parameter $\mathbf{z}$ for dynamic adjustment of ${\mathbf{y}_{m}}$ and ${\mathbf{y}_{p}}$. The two outputs are then combined and projected, followed by inverse shuffle and reshape operations to align with the input sequence. Finally, the output $\mathbf{F}_{l}^{m}$ of the module can be obtained through depth-wise convolution and feature flattening.
Testing with Monte Carlo averaging
Table 1: Quantitative comparison of pan-sharpening task on three datasets. Bold and underline show the best and second-best values, respectively. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
WorldView-II Gaofen-2 WorldView-III
[…]
INNformer AAAI’22 41.6903 0.9704 0.0227 0.9514 47.3528 0.9893 0.0102 0.5479 30.5365 0.9225 0.0747 3.1142
SFINet ECCV’22 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5479 30.5901 0.9236 0.0741 3.0798
MSDDN TGRS’23 41.8435 0.9711 0.0222 0.9478 47.4101 0.9895 0.0101 0.5414 30.8645 0.9258 0.0757 2.9581
PanFlowNet ICCV’23 41.8548 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 2.9531
FAME AAAI’24 42.0262 0.9723 0.0215 0.9172 47.6721 0.9898 0.0098 0.5542 30.9903 0.9287 0.0697 2.9531
DISPNet AAAI’24 41.8768 0.9702 0.0221 0.9157 47.4529 0.9898 0.0111 0.5532 30.0426 0.9153 0.0776 3.2620
Pan-mamba arxiv’24 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1740 0.9302 0.0698 2.8910
Ours - 42.3428 0.9734 0.0208 0.8840 47.9180 0.9903 0.0097 0.5109 31.4005 0.9327 0.0676 2.8098
Table 2: Evaluation of our method on real-world full-resolution scenes from the WV2 dataset. Bold and underline show the best and second-best values, respectively.
Metric SFIM Brovey GFPCA INNformer SFINet PanFlowNet FAME DISPNet Pan-mamba Ours
$D_{\lambda}$downarrow 0.1403 0.1026 0.1139 0.0995 0.1034 0.0966 0.0951 0.0944 0.0966 0.0941
$D_{S}$downarrow 0.1320 0.1409 0.1535 0.1305 0.1305 0.1274 0.1263 0.1264 0.1272 0.1266
QNR $\uparrow$ 0.7826 0.7728 0.7532 0.7858 0.7827 0.7910 0.7933 0.7938 0.7911 0.7939
Figure 5: Comparative visual experiments of several methods on WV3 datasets
We incorporate stochastic factors into the random shuffle operation, necessitating the marginalization of these factors in the generation of the final fusion. However, the random shuffle method presents a theoretical challenge due to the exponentially large number of potential models, making precise averaging of their predictions infeasible. Drawing inspiration from dropouts (Srivastava et al. 2014), we approximate the expected value of the entire model through layered expectations. Therefore, the computation of the random shuffle during testing can be expressed as follows, where $\mathbf{RM}=RandomMamba$:
$\mathbf{R}{\mathbf{M}^{test}}(x)=\underset{\mathbf{\mathcal{S}}}{\mathop{\mathbb{missing}}{E}}\,[\mathbf{RM}(x,\mathbf{\mathcal{S}})]$ (13)
In fact, estimating $\mathbf{R}{{\mathbf{M}}^{test}}$ based on the aforementioned equation requires enumerating all possible shuffle results, which imposes a significant computational burden. Therefore, we employ Monte Carlo averaging to estimate its expectation:
$\mathbf{R}{\mathbf{M}^{test}}(x)\approx\frac{1}{M}\sum\limits_{i=1}^{M}{\mathbf{RM}(x,{{\mathbf{\mathcal{S}}}_{i}})}$ (14)
Specifically, the input image is independently shuffled ${M}$ times, and then the ${M}$ outputs of $\mathbf{RM}$ are calculated. The mean of these outputs is computed to obtain the final estimate. When ${M}\to\infty$ is present, the Monte Carlo estimator closely approximates the true mean. Figure 4 illustrates the testing process we designed, which significantly reduces the actual testing time by incorporating multiple identical inputs in a mini-batch and utilizing GPUs for parallel computation.
Table 3: Quantitative comparison of MIF task on three datasets. bold and underline show the best and second-best values.
PET CT SPECT
[…]
PSLPT 0.888 0.548 0.373 0.815 0.675 0.502 0.432 0.810 0.850 0.359 0.325 0.933
EMFusion 0.943 0.685 0.783 1.221 1.190 0.552 0.475 1.266 0.885 0.665 0.692 1.212
MSRPAN 1.017 0.581 0.799 1.182 1.319 0.436 0.455 1.261 0.960 0.525 0.560 1.153
SwinFusion 1.642 0.703 0.683 0.725 1.537 0.522 0.545 0.579 1.678 0.744 0.720 0.684
Zero 0.950 0.635 0.774 1.162 0.768 0.320 0.582 1.199 1.046 0.582 0.681 1.180
U2Fusion 0.947 0.460 0.292 0.494 0.309 0.074 0.489 0.042 0.865 0.419 0.696 0.479
CDDFuse 1.481 0.650 0.765 1.227 1.589 0.526 0.530 1.224 0.995 0.786 0.719 1.169
Ours 1.491 0.797 0.741 1.256 1.580 0.546 0.592 1.330 1.470 0.820 0.747 1.240
Figure 6: Comparative visual experiments of several methods on MRI-PET datasets
Table 4: Ablation for our methods on three datasets. The PSNR/SSIM/SAM/ERGAS values on benchmarks are reported. The best results are shown in bold.
Ablation Variant WorldView-II Gaofen-2 WorldView-III
Baseline - 42.3428/0.9734/0.0208/0.8840 47.9180/0.9903/0.0097/0.5109 31.4005/ 0.9327/0.0676/2.8098
Core Operation RM w/o Random Shuffle 42.1460/0.9723/0.0211/0.9042 47.4284/0.9890/0.0102/0.5355 31.3915/0.9329/0.0676/2.8113
RCIM w/o Random Shuffle 42.2443/0.9729/0.0210/0.8919 47.6847/0.9899/0.0099/0.5292 31.3476/0.9325/0.0680/2.8230
RMIM w/o Random Shuffle 42.2136/0.9727/0.0211/0.8978 47.5787/0.9895/0.0100/0.5324 31.0971/0.9295/0.0704/2.9097
Loss Function
In accordance with established norms within this area, the loss function employed in our model for pan-sharpening is the L1 loss. In the MIF task, we use two input images and the fused image to calculate a composite loss function that includes L1 loss, SSIM loss, and gradient loss.
Experiments
Datasets and Benchmark
For the pan-sharpening task, we utilized datasets from WorldView-II (WV2), GaoFen2 (GF2), and WorldView-III (WV3), encompassing a variety of urban and natural scenes. We generated training samples following the Wald (Wald, Ranchin, and Mangolini 1997) protocol in the inaccessible ground truth. We conducted a thorough comparison of our proposed method against classical approaches and deep learning-based methods, including INNformer (Zhou et al. 2022a), SFINet (Zhou et al. 2022b), MSDDN (He et al. 2023), PanFlowNet (Yang et al. 2023), FAME (He et al. 2024b), DISPNet (Wang et al. 2024) and Pan-mamba (He et al. 2024a).
In the MIF task, we employed medical images from the Harvard Medical website, comprising pairs such as MRI-CT, MRI-PET, and MRI-SPECT. For this task, we compared our method with various deep learning-based multi-modal fusion techniques, including PSLPT (Wang, Deng, and Vivone 2024), EMFusion (Xu and Ma 2021), MSRPAN (Fu et al. 2021), SwinFusion (Ma et al. 2022), Zero (Lahoud and Süsstrunk 2019), U2Fusion (Xu et al. 2020), and CDDFuse (Zhao et al. 2023).
Implementation Details
All experiments were conducted using the PyTorch framework on two NVIDIA RTX 3060 GPUs. We trained for 500 epochs with a batch size of 4 for the pan-sharpening task and 200 epochs with a batch size of 1 for the medical image fusion (MIF) task. The network parameters were optimized using the Adam optimizer. The initial learning rate was set to $45e-$ 4, which was subsequently reduced to $85e-$ 8 using the CosineAnnealingLR scheduler over the specified epochs. For the pan-sharpening task, we randomly cropped training set images to obtain LRMS patches of size 32x32 and PAN images of size 128x128. For the MIF task, the training set images were cropped to 256x256.
Comparison with SOTA Methods
Pan-sharpening.
The experimental results on three datasets are presented in Table 1. Reference metrics, including PSNR, SSIM, SAM, and ERGAS (Alparone et al. 2007), are used to evaluate the fusion effect. The results demonstrate that the proposed method outperforms the SOTA methods across all metrics. Notably, in PSNR metric, our method achieves improvements of 0.1047, 0.2727, and 0.2301 dB over Pan-mamba, which has the second-best performance. Better PSNR and SSIM evidence that the fusion results transfers more information from the original image and experiences less distortion. In the qualitative comparison, Figure 5 shows a representative sample from the WV3 dataset. Our method performs better on the MSE graph, indicating that the fusion result is closer to the ground truth. Compared to other methods, our approach achieves a more accurate restoration of spectral and spatial details, highlighting the advantages of our fusion technique.
To further verify the generalization ability of our method in full-resolution scenes, we use three non-reference metrics, including ${{D}_{s}}$, ${{D}_{\lambda}}$ and QNR, to evaluate our method on the full WorldView-II dataset. Table 2 presents the experimental results. Our method outperforms the comparison methods across all results, demonstrating the strong adaptability of Shuffle Mamba in image fusion.
Medical Image Fusion.
The quantitative comparison results of four metrics on the MIF dataset are shown in Table 3. The proposed method performs well across almost all metrics, demonstrating its effectiveness in medical image fusion. In our experimental results, a higher VIF indicates closer alignment with human perception. Improved SCD, Qabf, and SSIM scores suggest that the fused image maintains higher similarity and experiences less distortion compared to the original images. Figure 6 presents a qualitative comparison of several methods on the MRI-PET dataset. Our method exhibits superior visual quality, a finding that is corroborated by the experimental metrics.
Ablation Experiments
To verify the effectiveness of the Random Shuffle operation, we conducted corresponding ablation experiments. The core operations of these experiments involved removing the shuffle operations from the three main modules. Notably, in the experiment where RM-related components were removed, the RM block in the RCIM module was excluded. The experimental results are presented in Table 4. Removing the shuffle operation resulted in a significant decrease in model performance, indicating that the new scanning strategy designed under the Shuffle Mamba framework effectively enhances the quality of multi-modal image fusion.
Due to Monte Carlo averaging, Shuffle Mamba makes trading memory and computational resources possible for improved performance. We studied the relationship between the number of samples, the PSNR index, and resource consumption. We experimented five times for each sample size, calculating the average and standard deviation of diverse metrics. Figure 7 (a) illustrates the PSNR trend, while Figure 7 (b) and (c) details the corresponding memory usage and the time required to process each image. As the number of samples increases, performance and resource consumption rise, enabling a trade-off between performance and efficiency. Additionally, Monte Carlo averaging enhances the theoretical robustness of the random shuffle operation, effectively improving the mean of PSNR while reducing the variance of the fusion result. More experimental results can be found in the supplementary materials.
Figure 7: The performance and costs in testing time and memory when choosing different numbers of samples.
Conclusion
In this paper, we replace the fixed scanning strategy used in current Mamba-based methods with a random shuffle-based scanning method and design a new Shuffle Mamba framework. This approach mitigates the bias introduced by fixed scanning methods and provides an global receptive field for multi-modal image fusion. During testing, we employ Monte Carlo averaging to account for the introduced random factors. Extensive experiments on two tasks demonstrate that our approach outperforms state-of-the-art methods and exhibits strong generalization capabilities.